How far apart must two charges of 4.2*10^4 C and -2.7*10^-4C be placed in order to have an electric potential energy with a magnitude of 2 J?
Answer #1

Here,

let the distance between the charges is x

for the electric potential energy

U = k * q1 * q2/x

2 = 9 *10^9 * 4.2 *10^4 * (2.7 *10^-4)/x

x = 5.103 *10^10 m

the distance between the charges is 5.103 *1010 m

if the charges are 4.2 *10^-4 and (2.7 *10^-4)

2 = 9 *10^9 * 4.2 *10^-4 * (2.7 *10^-4)/x

x = 510.3 m

the distance between the charges is 510.3 m